# Quantum metrology with a continuous-variable system

### v1

@@ F.1 General expression
$\Gamma^{-1}[\hat{\rho},\mathbf{\hat{r}}]$ $=\frac{\Omega\Gamma[\hat{\rho},\mathbf{\hat{r}}]\Omega^{T}}{\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]}.$ (195)
@@
$F_{Q}[\hat{\rho},\frac{1}{2}\hat{\bm{r}}^{T}H\hat{\bm{r}}]$ $=\frac{\,\mathrm{Tr}\{\Omega^{T}\frac{\Omega\Gamma[\hat{\rho},\mathbf{\hat{r}}]\Omega^{T}}{\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]}\Omega H\Gamma[\hat{\rho},\mathbf{\hat{r}}]H\}+\,\mathrm{Tr}\{\Omega H\Omega H\}}{1+(4\det\Gamma[\hat{\rho},\mathbf{\hat{r}}])^{-1}}+\langle\hat{\bm{r}}\rangle_{\hat{\rho}}^{T}H\Omega^{T}\frac{\Omega\Gamma[\hat{\rho},\mathbf{\hat{r}}]\Omega^{T}}{\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]}\Omega H\langle\hat{\bm{r}}\rangle_{\hat{\rho}}$
$=\frac{\,\mathrm{Tr}\{(\Gamma[\hat{\rho},\mathbf{\hat{r}}]H)^{2}\}+\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]\,\mathrm{Tr}\{(\Omega H)^{2}\}}{\frac{1}{4}+\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]}+\frac{\langle\hat{\bm{r}}\rangle_{\hat{\rho}}^{T}H\Gamma[\hat{\rho},\mathbf{\hat{r}}]H\langle\hat{\bm{r}}\rangle_{\hat{\rho}}}{\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]}.$ (196)
F.2 Gaussian rotation sensing
Of particular interest is the sensitivity under phase shifts generated by $\hat{n}=\hat{a}^{\dagger}\hat{a}=\frac{1}{2}(\hat{x}^{2}+\hat{p}^{2}-1)$. Up to an irrelevant constant, this corresponds to the quadratic Hamiltonian $\frac{1}{2}\hat{\bm{r}}^{T}H\hat{\bm{r}}$ with Hamiltonian matrix $H=I$. We obtain
$F_{Q}[\hat{\rho},\hat{n}]$ $=\frac{\,\mathrm{Tr}\{\Omega^{T}\Gamma^{-1}[\hat{\rho},\mathbf{\hat{r}}]\Omega\Gamma[\hat{\rho},\mathbf{\hat{r}}]\}-2}{1+(4\det\Gamma[\hat{\rho},\mathbf{\hat{r}}])^{-1}}+\langle\hat{\bm{r}}\rangle_{\hat{\rho}}^{T}\Omega^{T}\Gamma^{-1}[\hat{\rho},\mathbf{\hat{r}}]\Omega\langle\hat{\bm{r}}\rangle_{\hat{\rho}}$
$=\frac{\,\mathrm{Tr}\{\Gamma[\hat{\rho},\mathbf{\hat{r}}]^{2}\}-2\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]}{\frac{1}{4}+\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]}+\frac{\langle\hat{\bm{r}}\rangle_{\hat{\rho}}^{T}\Gamma[\hat{\rho},\mathbf{\hat{r}}]\langle\hat{\bm{r}}\rangle_{\hat{\rho}}}{\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]}$ (197)
@@ Appendix I Rotation sensing with pure Gaussian states
$\psi(x)=(2\pi\sigma_{x}^{2})^{-1/4}\exp\left(-\frac{\left(x-\sqrt{2}\alpha_{r}\right)^{2}}{4\sigma_{x}^{2}}+i\sqrt{2}\alpha_{i}\left(x-\sqrt{2}\alpha_{r}\right)+i\frac{\text{cov}_{xp}\left(x-\sqrt{2}\alpha_{r}\right)^{2}}{2\sigma_{x}^{2}}\right)\;,$ (206)

### v2

@@ F.2.2 Symplectic evolution with thermal noise
$\frac{\partial}{\partial\theta}\det\Gamma[\hat{\rho}(\theta,t),\mathbf{\hat{r}}$] $=e^{-\kappa t}(1-e^{-\kappa t})\left(n_{b}+\frac{1}{2}\right)\,\mathrm{Tr}{\left[\Omega(H\Gamma[\hat{\rho}(\theta,0),\mathbf{\hat{r}}]-\Gamma[\hat{\rho}(\theta,0),\mathbf{\hat{r}}]H)\right]},$ (211)
and
$\det\frac{\partial}{\partial\theta}\Gamma[\hat{\rho}(\theta,t),\mathbf{\hat{r}}$] $=e^{-2\kappa t}\det\frac{\partial}{\partial\theta}\Gamma[\hat{\rho}(\theta,0),\mathbf{\hat{r}}$]
$=e^{-2\kappa t}\left(4\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]\det H-\,\mathrm{Tr}{\left[\Gamma[\hat{\rho},\mathbf{\hat{r}}]H\right]}^{2}\right),$ (212)
where we used Eq. (206) for unitary evolutions. Inserting these expressions in Eq. (203) yields the QFI for Gaussian evolutions under the impact of thermal noise.
F.2.3 Displacement with thermal noise
Alternatively, we may consider the effect of thermal noise on a displacement, described by
$\mathcal{X}(t)$ $=e^{-\kappa t/2}\mathbb{I},$
$\mathcal{Y}(t)$ $=(1-e^{-\kappa t})\left(n_{b}+\frac{1}{2}\right)\mathbb{I},$
$\bm{d}(\theta)$ $=\theta\Omega\bm{u},$ (213)
As above, the parameter $\theta$ is proportional to the displacement amplitude with grows with time $t$, which, however, carries no information about $\theta$.
In this case, Eqs. (F.1) read
$\frac{\partial}{\partial\theta}\langle\hat{\bm{r}}\rangle_{\hat{\rho}(\theta,t)}$ $=\Omega\bm{u},$
$\Gamma[\hat{\rho}(\theta,t),\mathbf{\hat{r}}$] $=e^{-\kappa t}\Gamma[\hat{\rho},\mathbf{\hat{r}}]+(1-e^{-\kappa t})\left(n_{b}+\frac{1}{2}\right)\mathbb{I},$
$\frac{\partial}{\partial\theta}\Gamma[\hat{\rho}(\theta,t),\mathbf{\hat{r}}$] $=0.$ (214)
@@ F.2.3 Displacement with thermal noise
$\det\Gamma[\hat{\rho}(\theta,t),\mathbf{\hat{r}}$] $=e^{-2\kappa t}\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]+(1-e^{-\kappa t})^{2}\left(n_{b}+\frac{1}{2}\right)^{2}+e^{-\kappa t}(1-e^{-\kappa t})\left(n_{b}+\frac{1}{2}\right)\,\mathrm{Tr}{\left[\Gamma[\hat{\rho},\mathbf{\hat{r}}]\right]}$, (215)
$\frac{\partial}{\partial\theta}\det\Gamma[\hat{\rho}(\theta,t),\mathbf{\hat{r}}$] $=0,$ (216)
and
$\det\frac{\partial}{\partial\theta}\Gamma[\hat{\rho}(\theta,t),\mathbf{\hat{r}}$] $=0.$ (217)
From Eq. (203) we obtain
$F_{Q}[\hat{\rho}(\theta,t)]=\frac{\bm{u}^{T}\Gamma[\hat{\rho}(\theta,t),\mathbf{\hat{r}}]\bm{u}}{\det\Gamma[\hat{\rho}(\theta,t),\mathbf{\hat{r}}]}$. (218)
F.2.4 Rotation sensing: unitary evolution
Of particular interest is the sensitivity under phase shifts generated by $\hat{n}=\hat{a}^{\dagger}\hat{a}=\frac{1}{2}(\hat{x}^{2}+\hat{p}^{2}-1)$. Up to an irrelevant constant, this corresponds to the quadratic Hamiltonian $\frac{1}{2}\hat{\bm{r}}^{T}H\hat{\bm{r}}$ with Hamiltonian matrix $H=\mathbb{I}$. We obtain
$F_{Q}[\hat{\rho},\hat{n}$] $=\frac{4\,\mathrm{Tr}\left[\Gamma[\hat{\rho},\mathbf{\hat{r}}]\right]^{2}-16\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]}{4\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]+1}+\frac{\langle\hat{\bm{r}}\rangle_{\hat{\rho}}^{T}\Gamma[\hat{\rho},\mathbf{\hat{r}}]\langle\hat{\bm{r}}\rangle_{\hat{\rho}}}{\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]}$. (219)
@@ F.2.4 Rotation sensing: unitary evolution
F.2.5 Rotation sensing: Thermal noise
If the phase shift discussed above is affected by thermal noise, described by (F.2.2), the results of Sec. F.2.2 apply. Here, $H=\mathbb{I}$, and thus Eqs. (210) and (211) yield
$\det\Gamma[\hat{\rho}(\theta,t),\mathbf{\hat{r}}$] $=e^{-2\kappa t}\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]+(1-e^{-\kappa t})^{2}\left(n_{b}+\frac{1}{2}\right)^{2}+e^{-\kappa t}(1-e^{-\kappa t})\left(n_{b}+\frac{1}{2}\right)\,\mathrm{Tr}{\left[\Gamma[\hat{\rho},\mathbf{\hat{r}}]\right]}$, (222)
$\frac{\partial}{\partial\theta}\det\Gamma[\hat{\rho}(\theta,t),\mathbf{\hat{r}}$] $=0,$ (223)
respectively. The second result reflects the invariance of the state’s purity under changes of the parameter $\theta$ that only appears in the Hamiltonian part of the evolution. Hence, using $\det\Gamma[\hat{\rho}(\theta,t),\mathbf{\hat{r}}]=\det\Gamma[\hat{\rho}(0,t),\mathbf{\hat{r}}$], the resulting QFI is again independent of $\theta$ and reads
$F_{Q}[\hat{\rho}(0,t)]=e^{-2\kappa t}\frac{4\,\mathrm{Tr}\left[\Gamma[\hat{\rho},\mathbf{\hat{r}}]\right]^{2}-16\det\Gamma[\hat{\rho},\mathbf{\hat{r}}]}{4\det\Gamma[\hat{\rho}(0,t),\mathbf{\hat{r}}]+1}+e^{-2\kappa t}\frac{\langle\hat{\bm{r}}\rangle_{\hat{\rho}}^{T}\Gamma[\hat{\rho},\mathbf{\hat{r}}]\langle\hat{\bm{r}}\rangle_{\hat{\rho}}}{\det\Gamma[\hat{\rho}(0,t),\mathbf{\hat{r}}]}+e^{-\kappa t}(1-e^{-\kappa t})\left(n_{b}+\frac{1}{2}\right)\frac{\left|\langle\hat{\bm{r}}\rangle_{\hat{\rho}}\right|^{2}}{\det\Gamma[\hat{\rho}(0,t),\mathbf{\hat{r}}]}$. (224)
For $\kappa=0$ we recover the expression (219) of the unitary case.
Considering again a Gaussian state with covariance matrix (176) and first moments (220), we obtain
$F_{Q}[\hat{\rho}(0,t)]=\frac{4\alpha^{2}\left[(2n_{b}+1)\left(e^{\kappa t}-1\right)+(2n_{T}+1)(\cosh(2r)-\sinh(2r)\cos(2\varphi-\gamma))\right]}{2C-e^{2\kappa t}}+\frac{2(2n_{T}+1)^{2}\sinh^{2}(2r)}{C},$ (225)
where
$C=(2n_{b}+1)(2n_{T}+1)\cosh(2r)\left(e^{\kappa t}-1\right)+2n_{b}(n_{b}+1)\left(e^{\kappa t}-1\right)^{2}+2n_{T}(n_{T}+1)-e^{\kappa t}+e^{2\kappa t}+1.$ (226)
@@ Appendix I Rotation sensing with pure Gaussian states
$\psi(x)=(2\pi\sigma_{x}^{2})^{-1/4}\exp\left[-\frac{\left(x-\sqrt{2}\alpha_{r}\right)^{2}}{4\sigma_{x}^{2}}+i\sqrt{2}\alpha_{i}\left(x-\sqrt{2}\alpha_{r}\right)+i\frac{\text{cov}_{xp}\left(x-\sqrt{2}\alpha_{r}\right)^{2}}{2\sigma_{x}^{2}}\right]\;,$ (233)